OpenStudy (anonymous):

If anyone could help me with this Geometry, I would be super grateful. :) What is the lateral area of a rectangular prism if the base edges are 3 centimeters and 4 centimeters and the height is 6 centimeters? A. 84 cm² B. 92 cm² C. 150 cm² D. 220 cm² I'm honestly have no idea, but am open to learning if anyone would like to help me :)

OpenStudy (anonymous):

2h(w+l)=2*6*(3+4)=84m^2

OpenStudy (lgbasallote):

ahh prisms..ithe lateral surface are = perimeter of the base x height

OpenStudy (anonymous):

The lateral area is made up of four rectangular faces. (I'm assuming by "lateral" you're excluding the bases.) The area of a rectangle is found to be the product of its 2 dimensions. In this case, there are 2 height-width rectangles and two height-length rectangles. This can be expressed as: la=hw+hw+hl+hl

OpenStudy (anonymous):

Excellent! Thank you both so much, I was very much wondering which formula I should use, but now I know. Thank you both, you're a huge help. :)